# P.O. Box 15500, 00076 Aalto, Finland

### v1

@@ 5.2 Motivating Example of the Influence Spreading Model
We will now present an illustrative and motivating example of how the influence spreading algorithm works. This example will not only explain the influence spreading model, but it will also give insight into the results presented in Section 6.
Consider a simple four-node network presented in Figure 1, where edge weights denote influencing probabilities. While the network in Figure 1 is very simple, it shows some key features of the influence spreading model. Here we calculate the influence spreading probabilities for the maximum walk length of $L_{max}=5$. We consider two separate examples, first using node $1$ as the source and node $4$ as the target, and secondly using node $4$ as the source and node $1$ as the target. These calculations result in different end probabilities since there are more cycle possibilities when going from node $1$ to node $4$ than vice versa.
@@
Figure 1: A simple network where the edge weights are probabilities.
When node $1$ is the source and node $4$ is the target, we have in total ten different walks that have length 5 or less. These ten walks are illustrated in Figure 2. In contrast, when node $4$ is the source and node $1$ is the target, then we have in total only six different walks that have length 5 or less. These six walks are illustrated in Figure 3.
@@
Figure 2: All ten walks from node 1 to node 4 with walk length equal or less than 55. The numbers denote order of traversal along the edges of the network.
@@
Figure 3: All six walks from node 4 to node 1 with walk length equal to or less than 55. The numbers denote the order of traversal along the edges of the network.
When the influence spreading algorithm 1 is performed with node 4 as the source and node 1 as the target, we find six walks with the walk length shorter than or equal to $5$. The walks are then sorted in increasing lexicographic order as follows:
  1. 1.
    $4\rightarrow 3\rightarrow 1$
  2. 2.
    $4\rightarrow 3\rightarrow 2\rightarrow 1$
  3. 3.
    $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$
  4. 4.
    $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 1$
  5. 5.
    $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$
  6. 6.
    $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$.
The longest common prefix (LCP) walks in this lexicographic order are walks including $4\rightarrow 3\rightarrow 2\rightarrow 3$ and $4\rightarrow 3\rightarrow 4\rightarrow 3$. We can process walks with the same LCP walk lengths in any order. An example calculation of the combination of walk probabilities is combining the third $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$ and fifth $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$ walks. The third walk has probability $0.5^{4}$, the fifth walk has probability $0.5^{5}$, and their LCP walk has probability $0.5^{3}$. Therefore, the combined probability is $0.5^{4}+0.5^{5}-0.5^{4}\cdot 0.5^{5}/0.5^{3}=0.078125$. When the algorithm 1 is executed until completion, we get the probability of approximately $0.358$. This means that node 4 influences node 1 with an approximate probability of $0.358$. When the influence spreading algorithm 1 is performed with node 1 as the source and node 4 as the target, we get the result that node 1 influences node 4 with approximately $0.472$ probability. Hence, we see that node 1 has more influence on node 4 than vice versa. This result is explained by the different number of walk possibilities.
Figures 2 and 3 underscore a crucial aspect of the influence spreading model: Calculations are conducted only until the target node is reached in each walk. This feature exists because the influence spreading model focuses on the first arrival/influence. This feature results in a different number of possible walks when going from node 1 to node 4 and vice versa. There are four fewer paths from node 4 to 1 than there are from node 1 to 4. The position of node 1 gives it more possibilities for cycles before reaching the target node 4. In contrast, the position of node 4 allows for fewer cycle possibilities in the network before reaching target node 1. We also see that node 4 is on the periphery of this small network, while node 1 is more central. This is also an important observation when we discuss the results of other experiments in Section 6, where we will also refer to a more formalized definition of the periphery.
@@ 6 Results
In this section, we will compare the metrics defined by the influence spreading model to standard centrality metrics. Furthermore, we will further examine the behavior of the centrality metrics defined by the influence spreading model.
The influence spreading model defines its own betweenness centrality (7) and also new centrality metrics out-centrality (5) and in-centrality (6). The compared standard centrality metrics are betweenness centrality (3), degree centrality (1) and closeness centrality (2).
@@
We frequently characterize nodes as peripheral or central. Borgatti and Everett proposed formalizations for the concept of core/peripheral structure in networks [borgatti2000]. One of the formalizations they proposed was a discrete model consisting of a core and a periphery. The core was defined as a cohesive subgraph in which nodes are connected to each other in a maximal sense, and the periphery consists of nodes that are loosely connected to the core but do not share the maximal cohesion with the core. Yanchenko and Sengupta explain the core as consisting of nodes that are densely connected to each other, and some of them also have connections to peripheral nodes, whereas peripheral nodes are only sparsely connected to each other [yanchenko2023]. We use the term peripheral nodes in the sense and meaning given by Yanchenko and Sengupta, and central nodes are interpreted as belonging to the core.
In Sections 6.1 and 6.2, we will have a detailed individual node-level examination of the behavior of the influence spreading model and its centrality metrics on two small networks. Using small networks allows us to examine and explain the behavior of the influence spreading model and its centrality metrics with node-level detail. In Section 6.3, we examine the collective network-level behavior of the influence spreading model and its centrality metrics on larger $1000$ node random graphs/networks. Finally, in Section 6.4, we show a few counterexamples where some of the phenomena we have found do not appear. All experiments in Sections 6.1, 6.2, and 6.3 use the maximum allowed walk length $L_{max}=20$. In Section 7, the effect of using shorter or longer $L_{max}$ is discussed.
@@ 6.1 Detailed Node-Level Analysis of the Influence Spreading Model on the Krackhardt Kite Network
We will examine the behavior of the influence spreading model in the Krackhardt kite network [krackhardt1990]. The Krackhardt kite shown in Figure 4 is a useful example to compare the influence spreading model against standard centrality metrics, as it reveals the differences of the influence spreading model’s centrality measures compared to standard centrality measures. The Krackhardt kite network is well known for producing a different highest ranked node for each of the calculations of standard betweenness (node 8), closeness (nodes 6 and 7), and degree centrality (node 4) [krackhardt1990].
@@
Figure 4: Krackhardt kite. This network was originally presented in [krackhardt1990].
@@
In Figures 6 and 7, we see that peripheral nodes attain larger in-centrality values for high edge probabilities than more central nodes. The reason for this phenomenon is found in the example of Section 5.2. When a more central node is exerting influence towards a peripheral node, it has more cycle possibilities than a peripheral node due to its position in the network. In contrast, when a more peripheral node is exerting influence towards a central node, it has fewer cycle possibilities than a central node due to its position in the network. It is also important to emphasize that if a target node has a high degree, then all of the edges leading to the target node are only allowed as the last edge on a walk. The influence spreading model focuses on the first arrival (first influence) and therefore peripheral nodes are influenced more than central nodes when edge probabilities are high, since high edge probabilities allow the influence to reach a target node even for very long paths.
@@
6.2 Detailed Node-Level Analysis of the Influence Spreading Model on Another Small Network
In Figure 8, a small 10-node network is presented. The same network was used in [landherr2010] to compare standard centrality measures with each other. This small network has an almost chain-like structure, where nodes exist mainly on a chain with some exceptions. Applying the influence spreading model on this small network results in unique values for all the centrality measures defined by the influence spreading model when the common edge probabilities are not equal to zero or one.
12345678109
Figure 8: A small network with 10 nodes. This network was originally used in [landherr2010].
In Figure 9, ISM betweenness values and rankings are plotted, and the rankings are compared to standard betweenness values. It is notable that the rankings shown in the figure are similar to each other. However, there are some notable differences. The standard betweenness centrality assigns value 0 to nodes 1, 8, and 3. In contrast, ISM betweenness for node 3 is higher than for nodes 1 and 8 for all edge probabilities. Since the influence spreading model also accounts for walks and paths that are not the shortest path, this distinction makes sense. Node 3 is indeed between other nodes on certain walks and paths that are not the shortest paths. Therefore, the influence spreading model finds information that using only the shortest paths cannot.
Figure 9: Influence spreading model betweenness values and rankings alongside standard betweenness centrality rankings for the small network displayed in Figure 8. Node labels are marked above the lines.
In Figure 9, we see that for the edge probability value $1$, the ISM betweenness values show four separate and distinct values, and in some cases, multiple nodes have an ISM betweenness value that converges to the same value. Nodes 5 and 6 attain the highest ISM betweenness values for high edge probabilities and converge at the edge probability value $1$. When either node 5 or 6 and its incident edges are removed from the network, the network is separated into two components, one containing four nodes and the other containing five nodes. Removing node 4 and its incident edges results in the first component containing three nodes and the second component containing six nodes. Removing node 2 or 7 results in the first component that contains one node and the second component that contains eight nodes. Hence, nodes 2 and 7 have converging ISM betweenness values. Nodes 1, 3, 8, 9, and 10 have converging ISM betweenness values as well, since the removal of any of them and their incident edges results in the network still staying connected.
In Figure 9, there are also other notable phenomena. Node 5 is ranked as the second node when standard betweenness is used. Using ISM betweenness, node 5 attains this same rank only for higher edge probabilities. This can be interpreted such that for lower edge probabilities, the influence from and to surrounding node clusters on both sides (cluster of nodes 2, 3, and 4 and cluster of nodes 6, 7, 9, and 10) does not come through very well. However, for higher edge probabilities, the influence coming from and to node clusters on each side of it comes through with more intensity. The other notable phenomenon is the behavior of the nodes (3, 9, and 10) that are not positioned on the spine of the chain-like network structure. Initially, nodes 9 and 10 have lower ISM betweenness values than node 3, but as the edge probability value increases, this situation is eventually reversed for the edge probability value $0.9$.
In Figure 10, the out-centrality and in-centrality rankings are compared to the standard closeness centrality and degree centrality rankings. Figure 10 shows that there is a similarity in the rankings between out-centrality and standard closeness centrality. Initially, in-centrality also has similarity to standard closeness centrality since its values are initially the same as out-centrality. Many nodes of the network have the same degree and therefore some rankings are shared between nodes. Shared rankings complicate comparisons, but we still see how the highest degree nodes 2, 4, 6, and 7 also attain the highest four ranks of out-centrality values for all edge probabilities. Similarly, nodes 3, 5, 9, and 10 are jointly ranked fifth for degree centrality, and they attain out-centrality ranks ranging from fifth to eighth for all edge probabilities. The peripheral nodes 1 and 8 are at the bottom of the rankings for both out-centrality and degree centrality.
Figure 10: Influence spreading model out-centrality and in-centrality rankings compared to standard closeness and degree centrality rankings for the small network displayed in Figure 8. Node labels are marked above the lines.
In Figure 10, we see again the reversing in-centrality ranking behavior first seen for the Krackhardt kite network in Figure 6. Although out-centrality and in-centrality initially share the same rankings, as edge weight increases, the relationship begins to reverse. With the highest edge probabilities, the in-centrality rankings have fully reversed compared to out-centrality rankings.
Figure 10 shows us that initially, nodes on the periphery of the network (nodes 1 and 8) are influenced the least amount by other nodes simply because the influence of other nodes hardly reaches them due to low edge probabilities. When edge probabilities reach a sufficiently high level, the peripheral nodes become the most influenced nodes in the network.
On the other hand, Figure 10 shows that out-centrality rankings stay relatively stable for all edge probabilities. What this means is that the most influential nodes remain most influential regardless of the edge probability value. However, there is a notable phenomenon in the out-centrality rankings. Nodes 6, 7, and 8 switch places with nodes 4, 2, and 1, respectively, as the edge probability value increases. Nodes 4 and 6, nodes 2 and 7, and nodes 1 and 8 have similar positions in the network chain. Initially, at low edge probabilities, node 6 ranks higher than node 4, node 7 ranks higher than node 2, and node 8 ranks higher than node 1. One possible explanation for this is that nodes 6, 7, and 8 have greater closeness centrality compared to their counterparts—nodes 4, 2, and 1, respectively. This higher closeness centrality allows nodes to exert more influence, even at lower edge probabilities. Beginning at the edge probability of $0.5$, nodes 4, 2, and 1 start to surpass their counterpart nodes 6, 7, and 8, respectively.
Figure 10 also shows that for low edge probabilities, nodes on the periphery (nodes 1 and 8) receive relatively little influence from other nodes since their in-centrality rankings are low. We can infer that for low edge probabilities, the most influential nodes mainly influence nodes close to them, and their influence does not reach the periphery.
In Figure 11, we see the out-centrality and in-centrality values for the small network. It is notable that, unlike the Krackhardt kite network, out-centrality values have less dispersion. This reduced dispersion between values is due to the chain-like structure of this small network, which does not allow nodes to have wildly different positions in the network.
Figure 11: Influence spreading model out-centrality and in-centrality values for the small network displayed in Figure 8. The lowest probability is $0.01$ and thereafter probabilities increase from $0.1$ to $1.0$ in $0.1$ increments.
@@ 6.3 Network-Level Analysis of the Influence Spreading Model in Random Networks
In this section, we examine the behavior of the influence spreading model on larger randomly generated networks and compare the results to standard centrality metrics. We use scatter plots and Pearson correlation coefficients in our comparisons. The sample Pearson correlation coefficient $r_{xy}$ for a sequence of observation pairs $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ is defined as
@@
where $\bar{x}=(1/n)\sum_{i=1}^{n}x_{i}$ and $\bar{y}=(1/n)\sum_{i=1}^{n}y_{i}$ [kaptein2022].
@@
We have used stochastic network models to generate the random networks used in this section. The stochastic network models used are Erdős-Rényi random graph model [erdos1959], Watts-Strogatz small world model [watts1998] and Barabási-Albert preferential attachment model [barabasi1999]. Erdős-Rényi model comes from the mathematical tradition of random graph theory, whereas Watts-Strogatz and Barabási-Albert models are innovations of network science. These three random network models each capture a different type of network with their characteristic properties and provide an aggregate representation of many individual networks that share these properties. The characteristic properties of the random networks produced by these stochastic network models are very well known, and they have been extensively studied in the literature. Therefore, these three random network models provide varied, yet controlled conditions to assess the influence spreading model and its centrality metrics compared to standard centrality metrics.
@@
We also note that we performed the same or similar experiments on the well-known Les Misérables network [knuth1993] and on another random network generated by using the extension of the Barabási-Albert model proposed by Holme and Kim [holme2002]. However, the Les Misérables network data set is quite small with 77 nodes and 254 edges, while similar phenomena appeared in the results as in our chosen three random networks. Therefore, the inclusion of the Les Misérables network was considered redundant. Furthermore, the random network produced by the method proposed by Holme and Kim displayed results similar to Barabási-Albert random network. This is not entirely surprising, since the essential difference between these random network models is that the method of Holme and Kim can create high clustering, whereas the Barabási-Albert model cannot.
@@
The model widely known as the Erdős-Rényi model (or less widely known as the Erdős-Rényi-Gilbert model) has two formulations [erdos1959, gilbert1959]. In the first formulation, the number of nodes and edges is fixed, and then a graph is randomly selected out of all possible graphs. In the second formulation, the number of nodes is fixed, and edges are created between all pairs of nodes independently of all other edges with probability $p$. In this study, we use the second formulation and denote a random Erdős-Rényi network with $\text{ER}(n,p)$, where $n$ is the number of nodes, and $p$ is the independent edge creation probability.
The Erdős-Rényi model has been thoroughly studied and its properties are well known [newman2003]. The Erdős-Rényi model creates networks with low clustering, and its degree distribution has been shown to be the Poisson distribution.
@@
The Watts-Strogatz model uses a regular ring lattice of (fixed number of) $n$ nodes and $k$ neighbors for each node as the starting point, after which edges are rewired with probability $p$ [watts1998]. Thus, with probability $p$, a node is connected to a uniformly randomly selected non-neighbor (duplicate edges are not allowed) and with probability 1p1-p the edge remains unchanged. With this method of network construction, Watts and Strogatz aimed to develop a mixture of order and randomness to create more realistic networks. The parameter $p$ controls the amount of randomness. The Watts-Strogatz model has the ability to create high clustering and short average path lengths between nodes. Due to the short average path lengths, it is also called the small-world model. The exact degree distribution of the Watts-Strogatz model is derived in [albert2002]. When $p=0$, the degree distribution is constant $k$. When $p$ grows, the degree distribution begins to resemble the shape of the degree distribution of the Erdős-Rényi random graph [albert2002] i.e. the degree distribution becomes approximately Poisson distributed. We denote a network produced by the Watts-Strogatz model by $\text{WS}(n,k,p)$.
@@
The Barabási-Albert model addresses some of the limitations of Erdős-Rényi model and Watts-Strogatz model [barabasi1999]. First, the Barabási-Albert model does not assume a fixed number of nodes, and instead, the number of nodes is allowed to grow. Second, Barabási-Albert model does not create edges between nodes with equal probability since a new edge is more likely to be connected to a node that already has a high degree. This is called preferential attachment, where high-degree nodes are more likely to receive more neighbors when the network size grows. The Barabási-Albert model begins from an initial graph and then continues adding nodes using preferential attachment until the desired network size has been reached. We denote the Barabási-Albert network with $\text{BA}(n,m)$, where $n$ is the desired number of nodes, and $m$ is the number of edges to be attached from a new node to existing nodes. The degree distribution of the Barabási-Albert model is scale-free.
@@
Many real-world networks display high clustering and non-Poissonian degree distributions [newman2003]. Scale-free (i.e. power-law) degree distributions have been proposed to be a common phenomenon for many real networks [albert2002]. However, the prevalence of scale-free degree distributions in real networks has also been recently disputed in [broido2019], where it is actually found that scale-free degree distributions are rare and the log-normal distribution provides an adequate modeling fit instead of a power-law distribution. A short overview on the prevalence of scale-free degree distributions in real networks can be found in [holme2019]. In light of this differing research on what degree distributions characterize real networks, we argue that testing the behavior of the influence spreading model on random networks that display either Poissonian or scale-free degree distributions is especially justified.
@@
We created three separate random networks with $1000$ nodes each using the aforementioned Erdős-Rényi, Watts-Strogatz and Barabási-Albert models. All random networks were created using the Python package NetworkX  [SciPyProceedings_11]. The parameters were chosen such that each random network would have roughly the same number of edges. The random network $\text{ER}(1000,0.01)$ contains $1000$ nodes and $4910$ edges. The random network $\text{WS}(1000,10,0.5)$ has $1000$ nodes and $5000$ edges. The random network $\text{BA}(1000,5)$ contains $1000$ nodes and $4975$ edges. All three random networks are connected. The networks were created using seed $1729$.
@@
In Figure 12, we have compared ISM betweenness centrality to standard betweenness centrality for three random networks produced by the Erdős-Rényi, Watts-Strogatz, and Barabási-Albert models. Figure 12 reveals how, for lower edge probabilities, there is quite a strong positive correlation between the ISM betweenness and the standard betweenness. As the edge probability value increases, this correlation becomes weaker and weaker. Eventually, for the edge probability value $0.9$, we see that multiple levels are formed for the ISM betweenness values. This phenomenon is likely caused by differences in how the network breaks down into components when a node is removed.

### v2

@@ 3.2 Motivating Example of the Influence Spreading Model
We will now present an illustrative and motivating example of how the influence spreading algorithm works. This example will not only explain the influence spreading model, but also give insight into the results presented in Section 4.
Consider an undirected four-node network presented in Figure 1, where edge weights denote influencing probabilities. While the network in Figure 1 is small, it shows some key features of the influence spreading model. Here we calculate the influence spreading probabilities for the maximum walk length of $L_{max}=5$. We consider two separate examples, first using node $1$ as the source and node $4$ as the target, and secondly using node $4$ as the source and node $1$ as the target. These calculations result in different end probabilities since there are more cycle possibilities when going from node $1$ to node $4$ than vice versa.
@@
Figure 1: A small undirected network where the edge weights are probabilities.
When node $1$ is the source and node $4$ is the target, we have in total ten different walks that have length 5 or less. These ten walks are illustrated in Figure 3(a). In contrast, when node $4$ is the source and node $1$ is the target, then we have in total only six different walks that have length 5 or less. These six walks are illustrated in Figure 3(b).
@@
(a) All ten walks from node 1 to node 4 with walk length equal or less than 55.
@@
(b) All six walks from node 4 to node 1 with walk length equal to or less than 55.
Figure 3: Example walks with walk length equal or less than 55 for the network depicted in Figure 1. The numbers denote the order of traversal along the edges of the network.
When the influence spreading algorithm 1 is performed with node 4 as the source and node 1 as the target, we find six walks with the walk length shorter than or equal to $5$. The walks are then sorted in increasing lexicographic order as follows (i) $4\rightarrow 3\rightarrow 1$, (ii) $4\rightarrow 3\rightarrow 2\rightarrow 1$, (iii) $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$, (iv) $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 1$, (v) $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$, and (vi) $4\rightarrow 3\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$. The longest common prefix (LCP) walks in this lexicographic order are walks including $4\rightarrow 3\rightarrow 2\rightarrow 3$ and $4\rightarrow 3\rightarrow 4\rightarrow 3$. We can process walks with the same LCP walk lengths in any order. An example calculation of the combination of walk probabilities is combining the third $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 1$ and fifth $4\rightarrow 3\rightarrow 2\rightarrow 3\rightarrow 2\rightarrow 1$ walks. The third walk has probability $0.5^{4}$, the fifth walk has probability $0.5^{5}$, and their LCP walk has probability $0.5^{3}$. Therefore, the combined probability is $0.5^{4}+0.5^{5}-0.5^{4}\cdot 0.5^{5}/0.5^{3}=0.078125$. When the algorithm 1 is executed until completion, we get the probability of approximately $0.358$. This means that node 4 influences node 1 with an approximate probability of $0.358$. When the influence spreading algorithm 1 is performed with node 1 as the source and node 4 as the target, we get the result that node 1 influences node 4 with approximately $0.472$ probability. Hence, we see that node 1 has more influence on node 4 than vice versa. This result is explained by the different number of walk possibilities.
Figures 3(a) and 3(b) underscore a crucial aspect of the influence spreading model: The calculations are conducted only until the target node is reached on each walk. This feature exists because the influence spreading model focuses on the first arrival/influence. This feature results in a different number of possible walks when going from node 1 to node 4 and vice versa. There are four fewer paths from node 4 to 1 than there are from node 1 to 4. The position of node 1 gives it more possibilities for cycles before reaching the target node 4. In contrast, the position of node 4 allows for fewer cycle possibilities in the network before reaching target node 1. We also see that node 4 is on the periphery of this small network, while node 1 is more central. This is also an important observation when we discuss the results of other experiments in Section 4, where we will also refer to a more formalized definition of the periphery.
@@ 4 Results
In this section, we will compare the metrics defined by the influence spreading model to standard centrality metrics. Furthermore, we will further examine the behavior of the centrality metrics defined by the influence spreading model. The influence spreading model defines its own betweenness centrality (7) and also new centrality metrics, i.e., out-centrality (5) and in-centrality (6). The compared standard centrality metrics are betweenness centrality (3), degree centrality (1) and closeness centrality (2).
@@
We frequently characterize nodes as peripheral or central. Borgatti and Everett proposed formalizations for the concept of core/peripheral structure in networks [5]. One of the formalizations they proposed was a discrete model consisting of a core and a periphery. The core was defined as a cohesive subgraph in which nodes are connected to each other in a maximal sense, and the periphery consists of nodes that are loosely connected to the core but do not share the maximal cohesion with the core. Yanchenko and Sengupta explain the core as consisting of nodes that are densely connected to each other, and some of them also have connections to peripheral nodes, whereas peripheral nodes are only sparsely connected to each other [31]. We use the term peripheral nodes in the sense and meaning given by Yanchenko and Sengupta, and central nodes are interpreted as belonging to the core.
In Section 4.1, we will have a detailed individual node-level examination of the behavior of the influence spreading model and its centrality metrics on a small network. Using a small network allows us to examine and explain the behavior of the influence spreading model and its centrality metrics with node-level detail. In Section 4.2, we examine the collective network-level behavior of the influence spreading model and its centrality metrics on larger $1000$ node random graphs/networks. All experiments in Sections 4.1, and 4.2 use the maximum allowed walk length $L_{max}=20$. In Section 5, the effect of using shorter or longer $L_{max}$ is discussed.
@@ 4.1 Detailed Node-Level Analysis of the Influence Spreading Model on the Krackhardt Kite Network
We will examine the behavior of the influence spreading model in the Krackhardt kite network [20]. The Krackhardt kite shown in Figure 4 is a useful example to compare the influence spreading model against standard centrality metrics, as it reveals the differences of the influence spreading model’s centrality measures compared to standard centrality measures. The Krackhardt kite network is well known for producing a different highest ranked node for each of the calculations of standard betweenness (node 8), closeness (nodes 6 and 7), and degree centrality (node 4) [20].
@@
Figure 4: Krackhardt kite. This network was originally presented in [20].
@@
In Figures 6 and 7, we see that peripheral nodes attain larger in-centrality values for high edge probabilities than more central nodes. The reason for this phenomenon is found in the example of Section 3.2. When a more central node is exerting influence towards a peripheral node, it has more cycle possibilities than a peripheral node due to its position in the network. In contrast, when a more peripheral node is exerting influence towards a central node, it has fewer cycle possibilities than a central node due to its position in the network. It is also important to emphasize that if a target node has a high degree, then all of the edges leading to the target node are only allowed as the last edge on a walk. The influence spreading model focuses on the first arrival (first influence) and therefore peripheral nodes are influenced more than central nodes when edge probabilities are high, since high edge probabilities allow the influence to reach a target node even for very long walks.
@@ 4.2 Network-Level Analysis of the Influence Spreading Model in Random Networks
In this section, we examine the behavior of the influence spreading model on larger randomly generated networks and compare the results to standard centrality metrics. We use scatter plots, Pearson correlation coefficients and Spearman’s rank correlation coefficients in our comparisons. The sample Pearson correlation coefficient $r_{xy}$ for a sequence of observation pairs $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ is defined as
@@
where $\bar{x}=(1/n)\sum_{i=1}^{n}x_{i}$ and $\bar{y}=(1/n)\sum_{i=1}^{n}y_{i}$ [18]. The sample Pearson correlation coefficient is also called the product-moment correlation coefficient.
Spearman’s rank correlation coefficient calculates the product moment correlation coefficient between the ranks of pairs of observations [18]. If the data contains the same exact values, then ranks for those observations are tied, i.e. they have the same rank. For observations that are tied, Spearman’s rank correlation coefficient assigns the average of the untied ranks that would be assigned to these observations if no ties occurred. Formally, Spearman’s rank correlation coefficient $r_{S}$ is defined as
$r_{S}=\frac{\sum_{i=1}^{n}(R^{x}_{i}-\bar{R}^{x})(R^{y}_{i}-\bar{R}^{y})}{\sqrt{\left(\sum_{i=1}^{n}(R^{x}_{i}-\bar{R}^{x})^{2}\right)\left(\sum_{i=1}^{n}(R^{y}_{i}-\bar{R}^{y})^{2}\right)}},$
where $R^{x}_{i}$ and $R^{y}_{i}$ are the ranks of observations $x_{i}$ and $y_{i}$, respectively, $\bar{R}^{x}=(1/n)\sum_{i=1}^{n}R^{x}_{i}$ and $\bar{R}^{y}=(1/n)\sum_{i=1}^{n}R^{y}_{i}$. It should be noted that Spearman’s rank correlation coefficient does not change if the data is transformed in a monotonic way [18]. On the other hand, Pearson correlation coefficient measures the strength of the linear relationship between observations. Therefore, these correlation coefficients can quantify the relationship between observations in different ways.
@@
We have used stochastic network models to generate the random networks used in this section. The stochastic network models used are Erdős-Rényi random graph model [11], Watts-Strogatz small world model [30] and Barabási-Albert preferential attachment model [2]. Erdős-Rényi model comes from the mathematical tradition of random graph theory, whereas Watts-Strogatz and Barabási-Albert models are innovations of network science. These three random network models each capture a different type of network with their characteristic properties and provide an aggregate representation of many individual networks that share these properties. The characteristic properties of the random networks produced by these stochastic network models are very well known, and they have been extensively studied in the literature. Therefore, these three random network models provide varied, yet controlled conditions to assess the influence spreading model and its centrality metrics compared to standard centrality metrics.
@@
We also note that we performed the same or similar experiments on the well-known Les Misérables network [19] and on another random network generated by using the extension of the Barabási-Albert model proposed by Holme and Kim [16]. However, the Les Misérables network data set is quite small with 77 nodes and 254 edges, while similar phenomena appeared in the results as in our chosen three random networks. Therefore, the inclusion of the Les Misérables network was considered redundant. Furthermore, the random network produced by the method proposed by Holme and Kim displayed results similar to Barabási-Albert random network. This is not entirely surprising, since the essential difference between these random network models is that the method of Holme and Kim can create high clustering, whereas the Barabási-Albert model cannot.
@@
The model widely known as the Erdős-Rényi model (or less widely known as the Erdős-Rényi-Gilbert model) has two formulations [11, 14]. In the first formulation, the number of nodes and edges is fixed, and then a graph is randomly selected out of all possible graphs. In the second formulation, the number of nodes is fixed, and edges are created between all pairs of nodes independently of all other edges with probability $p$. In this study, we use the second formulation and denote a random Erdős-Rényi network with $\text{ER}(n,p)$, where $n$ is the number of nodes, and $p$ is the independent edge creation probability. The Erdős-Rényi model has been thoroughly studied and its properties are well known [25]. The Erdős-Rényi model creates networks with low clustering, and its degree distribution has been shown to be the Poisson distribution.
@@
The Watts-Strogatz model uses a regular ring lattice of (fixed number of) $n$ nodes and $k$ neighbors for each node as the starting point, after which edges are rewired with probability $p$ [30]. Thus, with probability $p$, a node is connected to a uniformly randomly selected non-neighbor (duplicate edges are not allowed) and with probability 1p1-p the edge remains unchanged. With this method of network construction, Watts and Strogatz aimed to develop a mixture of order and randomness to create more realistic networks. The parameter $p$ controls the amount of randomness. The Watts-Strogatz model has the ability to create high clustering and short average path lengths between nodes. Due to the short average path lengths, it is also called the small-world model. The exact degree distribution of the Watts-Strogatz model is derived in [1]. When $p=0$, the degree distribution is constant $k$. When $p$ grows, the degree distribution begins to resemble the shape of the degree distribution of the Erdős-Rényi random graph [1] i.e. the degree distribution becomes approximately Poisson distributed. We denote a network produced by the Watts-Strogatz model by $\text{WS}(n,k,p)$.
@@
The Barabási-Albert model addresses some of the limitations of Erdős-Rényi model and Watts-Strogatz model [2]. First, the Barabási-Albert model does not assume a fixed number of nodes, and instead, the number of nodes is allowed to grow. Second, Barabási-Albert model does not create edges between nodes with equal probability since a new edge is more likely to be connected to a node that already has a high degree. This is called preferential attachment, where high-degree nodes are more likely to receive more neighbors when the network size grows. The Barabási-Albert model begins from an initial graph and then continues adding nodes using preferential attachment until the desired network size has been reached. We denote the Barabási-Albert network with $\text{BA}(n,m)$, where $n$ is the desired number of nodes, and $m$ is the number of edges to be attached from a new node to existing nodes. The degree distribution of the Barabási-Albert model is scale-free.
@@
Many real-world networks display high clustering and non-Poissonian degree distributions [25]. Scale-free (i.e. power-law) degree distributions have been proposed to be a common phenomenon for many real networks [1]. However, the prevalence of scale-free degree distributions in real networks has also been recently disputed in [9], where it is actually found that scale-free degree distributions are rare and the log-normal distribution provides an adequate modeling fit instead of a power-law distribution. A short overview on the prevalence of scale-free degree distributions in real networks can be found in [17]. In light of this differing research on what degree distributions characterize real networks, we argue that testing the behavior of the influence spreading model on random networks that display either Poissonian or scale-free degree distributions is especially justified.
@@
We created three separate random networks with $1000$ nodes using the aforementioned Erdős-Rényi, Watts-Strogatz and Barabási-Albert models. All random networks were created using the Python package NetworkX (version 2.4) [15]. The parameters were chosen such that each random network would have roughly the same number of edges. The random network $\text{ER}(1000,0.01)$ contains $1000$ nodes and $4910$ edges. The random network $\text{WS}(1000,10,0.5)$ has $1000$ nodes and $5000$ edges. The random network $\text{BA}(1000,5)$ contains $1000$ nodes and $4975$ edges. All three random networks are connected, and were created using seed $1729$.
@@
In Figure 8, we have compared ISM betweenness centrality to standard betweenness centrality for three random networks produced by the Erdős-Rényi, Watts-Strogatz, and Barabási-Albert models. Figure 8 reveals how, for lower edge probabilities, there is a positive linear relationship between the ISM betweenness and the standard betweenness. As the edge probability value increases, this relationship becomes weaker and weaker. Eventually, for the edge probability value $0.9$, we see that multiple levels are formed for the ISM betweenness values. This phenomenon is likely caused by differences in how the network breaks down into components when a node is removed.